OpenStudy (anonymous):

dy/dx = (2y+1)/x how do you separate this???

OpenStudy (anonymous):

when taking partial derivative you treat anything but the x as a constant so your basically taking the derivative of 1/x which is = -1/x^2 so your answer is -(2y+1)/x^2

OpenStudy (anonymous):

noo for differential equations!

jimthompson5910 (jim_thompson5910):

dy/dx = (2y+1)/x dy/(2y+1) = dx/x integrate both sides with respect to each variable

OpenStudy (dan815):

ill tell you!